Question1 Soft Touch Company sells leather furniture. The following schedule relates to the companys inventory for the month of April Cost Sales Apri 1 Beginning inventory 73 units $43,654 30,233 3 Purchase 5 Sale 11 Purchase 15 Sale 22 Sale 28 Purchase 49 units 33 units 27 units 53 units 36 units 51 units $37,554 17,577 63,282 41,256 34,374 Soft Touch uses the periodic inventory system Your answer is partially correct. Try again Calculate Soft Touch Companys cost of goods sold, gross margin, and ending inventory using: i. FIFO ii. Weighted-average (Round calculations for cost per unit to 2 decimal places, e.g. 10.52. Round answers under weighted-average to 2 decimal places, e.g. 61,052.79.) Cost of Goods Sold Gross Margin Ending Invento 73887s 0.48 51951 i. FIFO 76755.08 0.46 49072.92 ii. Weighted-average
Answer #1

Cost of units available for sale = $43,654 + $30,233 + $17,577 + $34,374

= $125,838

Number of units available for sale = 73 + 49 + 27 + 51

= 200

Sales = $37,554 + $63,282 + $41,256

= $142,092

Units sold = 33 + 53 + 36

= 122

Units in Ending inventory = Units available for sale - Units sold

= 200 - 122

= 78

FIFO Method :

Under the First in first out (FIFO) method of inventory valuation, Cost of goods sold consists of the units from beginning inventory and earliest purchases. Ending inventory consists of the units from recent purchases.

Ending inventory of 78 units consists of 51 units from April 28 purchases and 27 units from April 11 purchases.

Ending inventory = $34,374 + $17,577

= $51,951

Cost of goods sold = Cost of units available for sale - Ending inventory

= $125,838 - $51,951

= $73,887

Gross Margin = Sales - Cost of goods sold

= $142,092 - $73,887

= $68,205

Weighted-average method :

Cost per unit = Cost of units available for sale / Number of units available for sale

= $125,838 / 200

= $629.19

Ending inventory = 78 units * $629.19 per unit

= $49,076.82

Cost of goods sold = Cost of units available for sale - Ending inventory

= $125,838 - $49,076.82

= $76,761.18

Gross Margin = Sales - Cost of goods sold

= $142,092 - $76,761.18

= $65,330.82